OpenStudy (anonymous):

A theater has 30 seats in the first row, 33 seats in the second row, 36 seats in the third row, and so on in the same increasing pattern. If the theater has 15 rows of seats, how many seats are in the theater?

OpenStudy (anonymous):

use a calculater first add 3 to 36 until you get 2 15 then add everything up

OpenStudy (anonymous):

39,42,45 and so on

OpenStudy (texaschic101):

common difference is 3. an = a1 + (n - 1) * d n = the term you are looking for....15 a1 = first term....30 d = common difference....3 now we sub a15 = 30 + (15 - 1) * 3 a15 = 30 + 14 * 3 a15 = 30 + 42 a15 = 72 That is not the total....we just had to find the 15th term to use in the sum formula. sn = n(a1 + an)/2 sn = sum of n terms a1 = first term ...30 an = 15th term...72 now we sub s15 = 15(30 + 72)/2 s15 = 15(102) / 2 s15 = 1530/2 s15 = 765 Total seats in the theater are 765

OpenStudy (texaschic101):

Learn how to do it here... http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm

OpenStudy (anonymous):

Graph a triangle (LMN) and reflect it over the line y=x to create triangle L'M'N'. Part 1: Describe the transformation using words. (3 points) Part 2: Draw a line segment from point L to the reflecting line, and then draw a line segment from point L' to the reflecting line. What do you notice about the two line segments you drew? (2 points) Part 3: Do you think you would see the same characteristic if you drew the line segment connecting M with the reflecting line and then M' with the reflecting line? How do you know? (3 points

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

can you help me with this ?